OpenStudy (anonymous):

Describe 3x - 4 + 4 = 11 + 4

OpenStudy (anonymous):

I guess this is how you do it... 3x-4+4=11+4 -4 and 4 would be canceled out because it would be 0 so that leaves with 3x=11+4 then 11+4 is 15 so thats 3x=15 divide 3 by 15 and x will equal 5